OpenStudy (anonymous):

...Easy Medal!! Simplify 4 to the 7th over 5 squared all raised to the 3rd power. 4 to the 10th over 5 to the 5th 4 to the 4th over 5 4 to the 21st over 5 to the 6th 12 to the 7th over 15 squared

OpenStudy (anonymous):

So is it like this? \[\large \sf \huge ( \large \frac{4^{7}}{5^{2}} \huge ) ^{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

Then you would use the exponent rule: \[\large \sf (b^{n})^m~=~b^{n \times m}\]

OpenStudy (anonymous):

which answer would that be

OpenStudy (anonymous):

\[\large \sf (\frac{b^{n}}{a^{l}})^m~=~\frac{b^{n \times m}}{a^{l \times m}}\]

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

That's not good

OpenStudy (anonymous):

..